[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Nov 27, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Nov 27 17:02:13 UTC 2017


HSU Shift Log for Mon Nov 27 2017 12:02:07 GMT-0500 (EST)

Shift: H

Forecaster: M Brennan

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: For the medium range hotline call, the phone didn't  
ring here at NHC when I dialed A1, but it did ring at WPC.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171127/80138165/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list